# All-weather Multi-Modality Image Fusion: Unified Framework and 100k Benchmark

### v2

@@ IV-D1 Fusion Performance Comparison
We selected 50 pairs of infrared and visible source images from the AWMM 100k dataset for each of the three types of adverse weather: rain, haze, and snow, as the test set. The ground truth corresponding to the degraded images was used as input for all comparison methods. Table II presents the results of all methods across nine objective metrics, with the first column indicating the source image types used by each method. The top three scores for each metric are highlighted. Notably, Table II shows that the proposed method, even when performing fusion under adverse weather conditions, outperforms comparison methods operating under ideal conditions. In rain, haze, and snow, the proposed teacher model ranks in the top three for eight or more metrics, while the student model ranks in the top three for , , and metrics, respectively.
@@
(2) Analyzing the and metrics, it is worth noting that AWFusion-S performs slightly worse in rain and haze scenes compared to snow on the metric. This discrepancy arises because rain and haze are continuously distributed degradations, while snow is more localized. As a result, snow is slightly easier to process than rain and haze. The proposed algorithm does not rank among the top three in the metric for haze scenes. This metric evaluates image quality based on texture, local structural changes, and noise sensitivity, suggesting that the performance of algorithm in extracting local texture under haze is somewhat weaker than in other conditions.
(3) For the metrics and , which are inspired by human perception, the proposed method performs exceptionally well. This aligns with the results shown in Figs. 5, 6 and 7 and Fig. 9, where the method effectively integrates multi-modality information while eliminating interference, ensuring that the fusion results are not distorted by adverse weather or enhanced in ways that conflict with human visual perception.
In summary, the performance of the proposed algorithm in quantitative comparisons further validates its practical value in handling severe weather conditions, demonstrating excellent fusion capabilities.
Source | Method | Pub. | Background | Car | Person | Bike | Curve | Car Stop | Cuardrail | Color cone | Bump | mIoU |
U2Fusion[2] | TPAMI 2020 | 98.39 | 92.88 | 73.25 | 73.03 | 52.82 | 73.75 | 86.2 | 40.31 | 83.2 | 74.87 | |
LRRNet[4] | TPAMI 2023 | 98.51 | 93.18 | 69.62 | 73.89 | 65.18 | 77.95 | 84.59 | 35.95 | 88.96 | 76.43 | |
MURF[38] | TPAMI 2023 | 99.13 | 95.56 | 83.54 | 86.65 | 76.27 | 88.21 | 93.48 | 76.25 | 90.59 | 87.74 | |
TIM[35] | TPAMI 2024 | 98.22 | 91.17 | 67.94 | 69.4 | 59.53 | 67.74 | 83.42 | 30.59 | 85.59 | 72.62 | |
CoCoNet[50] | IJCV 2023 | 98.3 | 92.03 | 69.31 | 70.66 | 57.72 | 73.62 | 87.49 | 29.44 | 84.9 | 73.72 | |
TarDal[3] | CVPR 2022 | 98.39 | 92.29 | 70 | 68.47 | 62.07 | 74.59 | 87.9 | 35.65 | 85.42 | 75.01 | |
IGNet[32] | MM 2023 | 99.12 | 95.61 | 83.09 | 87.22 | 74.97 | 87.96 | 93.1 | 76.13 | 90.9 | 87.6 | |
DRMF[51] | MM 2024 | 99.16 | 95.14 | 82.36 | 88.2 | 84.27 | 90.68 | 91.67 | 80.35 | 92.08 | 89.32 | |
MUFusion[52] | InfFus 2023 | 99.23 | 95.17 | 83.84 | 88.43 | 84.1 | 88.9 | 92.94 | 80.7 | 86.12 | 88.82 | |
GT | CrossFuse[53] | InfFus 2024 | 98.04 | 90.3 | 67.43 | 68.51 | 58.58 | 71.81 | 68.13 | 29.92 | 68.7 | 69.04 |
Rain | 99.32 | 96.49 | 84.77 | 88.76 | 86.35 | 90.45 | 93.91 | 78.45 | 91.82 | 90.04 | ||
Haze | 99.19 | 95.69 | 81.51 | 81.94 | 85.21 | 89.5 | 91.57 | 72.73 | 93.57 | 87.88 | ||
Snow | AWFusion-S | - | 99.25 | 95.92 | 81.24 | 86.4 | 86.77 | 90.36 | 93.74 | 78.04 | 92.32 | 89.34 |
Source | Method | Pub. | People | Car | Bus | Lamp | Motorcycle | Truck | mAP@0.5 | mAP@[0.5:0.95] |
U2Fusion[2] | TPAMI 2020 | 0.816 | 0.903 | 0.901 | 0.716 | 0.711 | 0.801 | 0.808 | 0.512 | |
LRRNet[4] | TPAMI 2023 | 0.785 | 0.909 | 0.912 | 0.764 | 0.721 | 0.808 | 0.817 | 0.484 | |
MURF[38] | TPAMI 2023 | 0.815 | 0.885 | 0.881 | 0.711 | 0.675 | 0.8 | 0.794 | 0.498 | |
TIM[35] | TPAMI 2024 | 0.731 | 0.898 | 0.878 | 0.657 | 0.656 | 0.769 | 0.765 | 0.48 | |
CoCoNet[50] | IJCV 2023 | 0.791 | 0.889 | 0.889 | 0.696 | 0.641 | 0.762 | 0.778 | 0.49 | |
TarDal[3] | CVPR 2022 | 0.792 | 0.877 | 0.857 | 0.585 | 0.62 | 0.713 | 0.741 | 0.462 | |
IGNet[32] | MM 2023 | 0.793 | 0.859 | 0.836 | 0.532 | 0.554 | 0.721 | 0.716 | 0.439 | |
DRMF[51] | MM 2024 | 0.699 | 0.897 | 0.896 | 0.787 | 0.697 | 0.784 | 0.793 | 0.494 | |
MUFusion[52] | InfFus 2023 | 0.803 | 0.906 | 0.908 | 0.734 | 0.694 | 0.802 | 0.808 | 0.507 | |
CrossFuse[53] | InfFus 2024 | 0.692 | 0.891 | 0.871 | 0.671 | 0.674 | 0.771 | 0.762 | 0.473 | |
GT | AWFusion-S | - | 0.791 | 0.922 | 0.93 | 0.804 | 0.726 | 0.821 | 0.832 | 0.534 |
IV-D2 Semantic Segmentation Performance Comparison
@@
The proposed all-weather MMIF algorithms are designed to support real-world applications, such as autonomous driving, surveillance, and UAV navigation. In these scenarios, the goal is not just to generate visually enhanced images but to enable effective decision-making and analysis (e.g., object detection, semantic segmentation) from the fused images. To comprehensively evaluate the fusion performance under three types of severe weather, we introduced a downstream task. In Table III, we compare the performance of various algorithms using BANet [61] as the segmentation network. Unlike comparison methods, which receive clean source images as input, the proposed algorithm performs fusion under severe weather conditions. As shown in Table III, the proposed algorithm ranks first in mean intersection-over-union (mIoU) for rain scenes and second for snow scenes. Although the mIoU score for the haze scenario is not in the top three, it still secures fifth place. These experiments demonstrate that the proposed algorithm retains high-level semantic information in scenes with interference and accurately recognizes key targets such as roads, pedestrians, and vehicles.
IV-D3 Object Detection Performance Comparison
@@
In real-world applications, systems must adapt to varying weather and environmental conditions. Although our method is tailored for severe weather, it also delivers high-quality fusion results in normal conditions, such as clear, haze-free, and rain-free environments. This demonstrates not only its effectiveness in harsh weather but also its strong generalization ability in clean settings. In Table IV, we evaluated the object detection accuracies of all methods on the M3FD dataset using YOLOv7 [10]. As shown in Table IV, the proposed algorithm achieves the highest detection accuracy across all five categories, with both mAP@0.5 and mAP@[0.5:0.95] also obtaining the top scores. These results confirm that the proposed algorithm retains sufficient information and detail for accurate localization and identification in object detection. Moreover, it enhances detection accuracy by preserving and leveraging complementary information.
IV-E Extension to Multiple Adverse Weather Disturbances
@@
Compound degradation increases scene complexity, as rain, haze, and snow exhibit distinct physical characteristics: rain and snow cause localized occlusion, while haze reduces overall contrast. When combined, these effects create more challenging degradation. Although specific patterns can address individual disturbances, the model must exhibit strong adaptability and generalization to handle multiple disturbances simultaneously. To evaluate the generalization ability of the proposed algorithm under severe weather conditions, we conducted extended experiments with composite degradation. We used DT [62] as a preprocessing model to remove interference from visible images. In the comparison method, fusion is performed on the preprocessed image, while the proposed algorithm directly processes the fusion result on the original source image. As shown in Fig. 10, when two disturbances coexist in the source image, DT can only address one at a time, despite its effectiveness handle individual interferences. The residual degradation is carried over to the fusion result, leading to suboptimal outcomes. In contrast, the proposed method reconstructs high-fidelity scene details under complex weather conditions, preserving key features while avoiding issues like insufficient brightness or information loss.
IV-F Ablation Experiment
@@
We performed an ablation analysis on the six modules to verify their individual roles and effectiveness. The modules were grouped into four experimental sets, with the results visualized in Fig. 11 and quantitatively compared in Table V.
IV-F1 Effectiveness Analysis of the Fusion Module
@@ IV-F2 Effectiveness Analysis of the PFB
IV-F3 Effectiveness Analysis of the IFB and RFB
To improve the ability of network to estimate light transmission, we designed the IFB and RFB modules based on retina theory and constrained feature generation by reconstructing the loss. As shown in the third row of Fig. 11, removing these modules results in reduced lighting in the fusion output, likely due to inaccurate transmission map estimation and insufficient luminance compensation. This makes the absence of light recovery more noticeable. Additionally, Fig. 4 shows a significant drop in the , , and metrics.
IV-F4 Effectiveness Analysis of the Fusion Strategies
To leverage the low-rank and sparse features decomposed by the LLRR block, we designed the SFB to fuse sparse components and the STB to integrate low-rank components through context extraction. An ablation study was performed by removing the SFB and STB, replacing them with a maximum-value fusion strategy. As more feature channels are extracted from the source image, the simple fusion strategy impedes effective cross-modal interaction, restricting the extraction of salient and complementary features. This causes pixel redundancy in the fused output, reflected in the excessively high metric in Table V. The enlarged region in Fig. 11 further highlights increased noise in ground, complicating object recognition.
V Conclusion
In this study, we propose an all-weather MMIF method, focusing on the relationship between multi-modality image fusion and image restoration. Our framework shifts the network’s objective from pixel-perfect recovery to maximizing the representation of key scene information. The end-to-end design comprises two main modules: a fusion module and a restoration module. The fusion module decouples features from multiple modalities into low-rank and sparse components, which are merged interactively using specialized fusion strategies. Guided by a physical model, the restoration module restores clear details from blur pixels, ensuring high-quality fusion results. Additionally, we introduce a large-scale benchmark dataset, AWMM-100k, tailored for multi-modality image processing under adverse weather conditions. Experimental results confirm that the AWFusion surpasses existing methods in both adverse and normal conditions, with outstanding performance in downstream tasks, demonstrating its strong practical value.

### v3

@@ 4.4.1 Fusion Performance Comparison
We selected 50 pairs of infrared and visible source images from the AWMM 100k dataset for each of the three types of adverse weather: rain, haze, and snow, as the test set. The ground truth corresponding to the degraded images was used as input for all comparison methods. Table 2 presents the results of all methods across nine objective metrics, with the first column indicating the source image types used by each method. The top three scores for each metric are highlighted. Notably, Table 2 shows that the proposed method, even when performing fusion under adverse weather conditions, outperforms comparison methods operating under ideal conditions. In rain, haze, and snow, the proposed teacher model ranked in the top three across more than eight evaluation metrics. Meanwhile, the student model achieved top-three rankings on seven, four, and seven metrics under the respective conditions.
@@
(2) Analyzing the and metrics, it is worth noting that AWFusion-S performs slightly worse in rain and haze scenes compared to snow on the metric. This discrepancy arises because rain and haze are continuously distributed degradations, while snow is more localized. As a result, snow is slightly easier to process than rain and haze. The proposed algorithm is not among the top three in the metric for haze scenes. This metric evaluates image quality based on texture and local structural changes, suggesting that the performance of algorithm in extracting local texture under haze is somewhat weaker than in other conditions.
4.4.2 Semantic Segmentation Comparison
@@
The proposed all-weather MMIF algorithms are designed to support real-world applications, such as autonomous driving, surveillance, and UAV navigation. In these scenarios, the goal is not just to generate visually enhanced images but to enable effective decision-making and analysis (e.g., object detection, semantic segmentation) from the fused images. To comprehensively evaluate the fusion performance under three types of severe weather, we introduced a downstream task. In Table 3, we compare the performance of various algorithms using BANet [41] as the segmentation network. Unlike comparison methods, which receive clean source images as input, the proposed algorithm performs fusion under severe weather conditions. As shown in Table 3, the proposed algorithm ranks first in mean intersection-over-union (mIoU) for rain scenes and second for snow scenes. Although the mIoU score for the haze scenario is not in the top three, it still secures fifth place. These experiments demonstrate that the proposed algorithm retains high-level semantic information in scenes with interference and accurately recognizes key targets such as roads, pedestrians, and vehicles.
Methods Pub. Restor. ↑ ↑ ↑ ↑ ↓ ↑ ↑ ↑ ↑ Snow and Haze ControlFusion [48] NIPS 25 MODEM [57] 0.3408 0.3230 0.7097 0.4499 383.7061 0.3097 1.0466 0.4581 7.1471 GIFNet [7] CVPR 25 0.2950 0.2811 0.6826 0.4579 717.2248 0.3304 1.2148 0.3595 6.6378 SAGE [60] CVPR 25 0.3212 0.3630 0.7247 0.4445 443.1403 0.3259 1.1365 0.4524 7.0594 AWFusion-S _ _ 0.3578 0.3590 0.7494 0.4584 519.5943 0.3698 1.3985 0.4656 7.1554 Rain and Haze ControlFusion [48] NIPS 25 MODEM [57] 0.3385 0.3453 0.7036 0.4350 385.4859 0.2995 0.9922 0.4680 7.3067 GIFNet [7] CVPR 25 0.3023 0.2956 0.6924 0.4529 662.5418 0.3292 1.2082 0.3685 6.6517 SAGE [60] CVPR 25 0.3182 0.3880 0.7221 0.4296 436.8800 0.3138 1.0881 0.4516 7.0836 AWFusion-S _ _ 0.3775 0.3999 0.7530 0.4454 549.1318 0.3722 1.4212 0.4919 7.1886 Rain and Snow ControlFusion [48] NIPS 25 MODEM [57] 0.3551 0.3203 0.7258 0.4594 537.3284 0.3155 1.2346 0.4869 7.0407 GIFNet [7] CVPR 25 0.2924 0.2826 0.6588 0.4622 1121.0535 0.3113 1.1707 0.3739 6.6209 SAGE [60] CVPR 25 0.3138 0.3781 0.7579 0.4419 628.5957 0.3124 1.2184 0.4821 6.7914 AWFusion-S _ _ 0.3382 0.3622 0.6955 0.4699 258.0324 0.3198 1.3459 0.4910 7.1263
4.4.3 Object Detection Performance Comparison
@@
In real-world applications, systems must adapt to varying weather and environmental conditions. Although our method is tailored for severe weather, it also delivers high-quality fusion results in normal conditions, such as clear, haze-free, and rain-free environments. This demonstrates not only its effectiveness in harsh weather but also its strong generalization ability in clean settings. In Table 4, we evaluated the object detection accuracies of all methods on the M3FD dataset using YOLOv7 [55]. As shown in Table 4, the proposed algorithm achieves the highest detection accuracy across all five categories, with both mAP@0.5 and mAP@[0.5:0.95] also obtaining the top scores. These results confirm that the proposed algorithm retains sufficient information and detail for accurate localization and identification in object detection. Moreover, it enhances detection accuracy by preserving and leveraging complementary information.
4.5 Extension to Comparisons with “Restoration + Fusion” pipelines
To further validate the effectiveness of our method, we extend our experiments to comparisons with “Restoration + Fusion” pipelines. Specifically, we first apply image restoration algorithms to recover the degraded visible images, and then feed the restored visible images together with clean infrared images into SOTA IVIF methods. In contrast, our model directly takes degraded multi-modal images as input. We use two restoration methods, AdaIR [9] and MODEM [57], and compare our method with five advanced IVIF algorithms: Text-DiFuse [69], ControlFusion [48], GIFNet [7], SAGE [60], and OmniFuse [70]. Qualitative results are shown in Fig. 12, while quantitative results are reported in Table 5 and Table 6. For fairness, both restoration algorithms were retrained on the AWMM-100K dataset.
From Fig. 12, we can observe that although restoration algorithms can remove most degradation, the remaining artifacts and detail loss are inevitably passed to the subsequent fusion stage. In rainy scenes, the fusion results of Text-DiFuse, ControlFusion, and OmniFuse show varying degrees of detail loss, leading to insufficient scene contrast. GIFNet carries noticeable artifacts produced during restoration, which may interfere with downstream tasks such as target recognition and scene understanding. In hazy scenes, similar issues persist. GIFNet attempts to compensate for the lost details by overly enhancing infrared information, but this causes the final image to deviate from human visual perception. Since Text-DiFuse and OmniFuse cannot perceive degradation cues, they amplify the residual degradation from the restoration stage, further reducing image quality. In snowy scenes, snowflakes not only occlude structural details but also introduce effects similar to overexposure. As a result, GIFNet and SAGE fail to preserve scene brightness, leading to incomplete depth information. In contrast, the proposed method consistently delivers robust fusion performance across all three adverse conditions. It preserves fine details, maintains color fidelity, and effectively removes degradation, outperforming all “Restoration + Fusion” pipelines.
Table 5 and Table 6 present the quantitative comparisons using AdaIR-restored and MODEM-restored visible images. As reflected in the qualitative analysis, Text-DiFuse and OmniFuse obtain relatively low scores due to their significant loss of details. The proposed method surpasses all competing approaches across the three weather conditions and achieves the highest scores on most evaluation metrics. Overall, the proposed method outperforms the “Restoration + Fusion” pipelines in both qualitative and quantitative comparisons, further demonstrating its advantages for real-world applications.
4.6 Extension to Multiple Weather Disturbances
@@
Compound degradation increases scene complexity, as rain, haze, and snow exhibit distinct physical characteristics: rain and snow cause localized occlusion, while haze reduces overall contrast. When combined, these effects create more challenging degradation. Although specific patterns can address individual disturbances, the model must exhibit strong adaptability and generalization to handle multiple disturbances simultaneously. To evaluate the generalization ability of the proposed algorithm under severe weather conditions, we conducted extended experiments with composite degradation. In the comparison methods, the fusion operation is performed on pre-processed images. We adopt MODEM, a state-of-the-art adverse-weather removal model, to eliminate degradations in the visible images before fusion. Since MODEM is trained on AWMM-100K, it is capable of identifying and removing all three types of degradations. In contrast, our method directly generates fused results from the original degraded multi-modal inputs, without relying on any pre-processing.
As shown in Fig. 13, when two types of degradations appear simultaneously in the source images, the degradation removal capability of MODEM is significantly reduced. The residual degradation is carried over to the fusion result, leading to suboptimal outcomes. In contrast, the proposed method reconstructs high-fidelity scene details under complex weather conditions, preserving key features while avoiding issues like information loss. Table 7 presents the quantitative results of different methods under multiple weather degradations. As shown, the proposed method achieves the highest scores on six or more metrics across all three types of composite degradation.
4.7 Ablation Experiment
@@
We performed an ablation analysis on the six modules to verify their effectiveness. The modules were grouped into four experimental sets, with the results visualized in Fig. 14 and quantitatively compared in Table 8.
4.7.1 Effectiveness Analysis of the Fusion Module
@@
4.7.2 Effectiveness Analysis of the PFB
4.7.3 Effectiveness Analysis of the IFB and RFB
To improve the ability of network to estimate light transmission, we designed the IFB and RFB modules based on retina theory and constrained feature generation by reconstructing the loss. As shown in the third row of Fig. 13, removing these modules results in reduced lighting in the fusion output, likely due to inaccurate transmission map estimation and insufficient luminance compensation. This makes the absence of light recovery more noticeable. Additionally, Table 8 shows a significant drop in the , , and metrics.
4.7.4 Effectiveness Analysis of the Fusion Strategies
To leverage the low-rank and sparse features decomposed by the LLRR block, we designed the SFB to fuse sparse components and the STB to integrate low-rank components through context extraction. An ablation study was performed by removing the SFB and STB, replacing them with a maximum-value fusion strategy. As more feature channels are extracted from the source image, the simple fusion strategy impedes effective cross-modal interaction, restricting the extraction of salient and complementary features. This causes pixel redundancy in the fused output, reflected in the excessively high metric in Table 8. The enlarged region in Fig. 13 further highlights the increased noise in the ground, which complicates object recognition.
5 Discussion
Limitations and Future Work: The proposed algorithm offers an effective solution for IVIF under adverse weather conditions. While it demonstrates strong fusion performance, our experiments reveal that its computational efficiency leaves room for further improvement. Specifically, for an input of size 256×256, the algorithm requires 193 ms of inference time, with 889 GFLOPs and 19.359 million parameters. In future work, we aim to optimize the computational efficiency of the model, thereby enhancing its applicability in real-time and resource-constrained scenarios.
Conclusion: In this study, we propose an all-weather MMIF method, focusing on the relationship between multi-modality image fusion and image restoration. Our framework shifts the goal of the network from pixel-perfect recovery to maximizing the representation of key information from the scene. The end-to-end design comprises two main modules: a fusion module that decouples multi-modal features into low-rank and sparse components for interactive merging, and a restoration module guided by physical model to restore clear details from blur pixels, ensuring high-quality results. Additionally, we introduce a large-scale benchmark dataset, AWMM-100k, tailored for multi-modality image processing under adverse weather conditions. Experimental results confirm that the AWFusion surpasses existing methods in both adverse and normal conditions, with outstanding performance in downstream tasks, demonstrating its strong practical value.
Acknowledgement
This research was supported by the National Natural Science Foundation of China (No. 62201149), the Natural Science Foundation of Guangdong Province (No. 2024A1515011880), the Basic and Applied Basic Research of Guangdong Province (No. 2023A1515140077), the Research Fund of Guangdong-HongKong-Macao Joint Laboratory for Intelligent Micro-Nano Optoelectronic Technology (No. 2020B1212030010).